OpenStudy (anonymous):

Find the axis of symmetry of the graph of the function. f(x) = 2x2 - 16x + 27 A. x = -3 B. x = 4 C. x = -5 D. x = 6

OpenStudy (anonymous):

i keep getting an answer thats not one of the choices :( please help..

OpenStudy (anonymous):

-b/2a x==4

OpenStudy (anonymous):

howd u find that answer?

OpenStudy (anonymous):

i just found the vertex and the equation is -b/2a so its -(-16) / 2(2) = 4

OpenStudy (anonymous):

but what happens to the 27?

OpenStudy (anonymous):

If you complete the square, you can find the low point of the parabola. It's a low point since the parabola opens upwards.

OpenStudy (anonymous):

27 is just an intercept it is not needed

OpenStudy (anonymous):

oohhhok, so the axis of symmetry would be 4?

OpenStudy (anonymous):

2x2 - 16x + 27 = 2(x-4)^2 - 5

OpenStudy (anonymous):

So, low point at (4, -5) and the parabola opens up on either side symmetrically. So axis of symmetry at x=4.

OpenStudy (anonymous):

oohhok! thankyou!!